# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

### v1

@@ Abstract
Improving diesel engine efficiency and emission reduction have been critical research topics over several decades. Recent government regulations have shifted this focus to another important area related to engine health and performance monitoring. Although the advancements in the use of deep learning methods for system monitoring have shown promising results in this direction, designing efficient methods suitable for field systems remains an open research challenge. To this end, the objective of this study is to develop a computationally efficient neural network-based approach for identifying unknown parameters of a “mean value” diesel engine model to facilitate physics-based health monitoring and maintenance forecasting. We propose a hybrid method combining physics-informed neural networks (PINNs) and a deep neural operator (DeepONet) to predict unknown parameters and gas flow dynamics in a diesel engine. The operator network predicts independent actuator dynamics learnt through offline training, thereby reducing the PINN’s online computational cost. Additionally, empirical relationships in the mean-value model are represented using deep neural networks trained offline. To address PINN’s need for retraining with changing input scenarios (condition of the ODE/PDE or measured data), we propose two transfer learning (TL) strategies. The first strategy involves multi-stage transfer learning for parameter identification. While this method is computationally efficient as compared to online PINN training, improvements are required to meet field requirements. The second TL strategy focuses solely on training the output weights and biases of a sub-set of multi-head networks pre-trained on a larger dataset, substantially reducing computation time during online prediction. We also evaluate our model for epistemic and aleatoric uncertainty by incorporating dropout in pre-trained networks and Gaussian noise in the training dataset. This strategy offers a tailored, computationally inexpensive, and physics-based approach for parameter identification in diesel engine sub-systems for health monitoring and diagnosis.
Keywords DeepONet  $\cdot$ Inverse problem  $\cdot$ Transfer learning  $\cdot$ Few shot learning  $\cdot$ Digital twin  $\cdot$ System-of-Systems
@@ 1 Introduction
The health monitoring of internal combustion engines has evolved significantly over the past decades, driven by the need for improved performance, reliability, and maintenance efficiency. Historically, manufacturers developed various diagnostics and service tools to meet customer, operational, and service needs. However, traditional monitoring approaches often face challenges such as lack of fault isolation, limited prognostic capabilities, and robustness issues due to system variability and gradual degradation. For instance, system-level monitors may indicate a fault due to multiple slightly degraded components, none of which are individually degraded enough to trigger a fault on their own. Such scenarios frequently lead to trial-and-error methods for fault resolution and multiple service visits for end-users. Early strategies to address these challenges included the implementation of component-level monitoring with very tight thresholds. However, this approach proved problematic due to the increased risk of false positives and the challenges associated with developing and tuning high-fidelity monitors for each component. Consequently, various alternative solutions have been explored, including the use of complex logic algorithms, multiple degradation thresholds, and out-of-mission system tests. The proliferation of numerous small diagnostic components designed to cover all aspects of system monitoring has resulted in a highly complex and costly landscape for development, tuning, documentation, and risk management.
The present study aims to streamline the development of health monitoring systems, improve fault isolation, system health prognostics, and predictive maintenance capabilities. Our focus is on advancing system-level monitoring solutions that utilize all available data to provide high-fidelity monitoring of system performance while also delivering robust fault isolation and prognostic capabilities. While the addition of electronics, sensors, and control systems helps in improving system performance, timely health monitoring of the engine’s sub-systems is crucial to ensure optimal performance under all operating conditions. The integration of a wide array of sensors and actuators has enabled the assimilation of accurate data during engine operation. The availability of sensory data has catalyzed active research in the use of neural networks for engine modeling and fault diagnosis. The complex non-linear mapping learned by neural networks offers significant advantages for practitioners, given the inherent challenges in creating accurate physics-based models for combustion engines. By evaluating the important design parameters of the engine from field data recorded by sensors, it is possible to infer the health of the engine by comparing them with the actual design parameters. Numerous studies have demonstrated the application of neural networks in modelling different aspects of diesel engine [1, 2, 3, 4, 5, 6, 7, 8, 9]. However, majority of these works focus on data-driven models that rely on data measurements to predict engine states. In an effort to integrate known physics with measurement data, [10] proposed a hybrid model for predicting gas exchange processes in a diesel engine for individual sub-systems such as intake and exhaust manifold. This work focuses primarily on estimating the gas flow states for individual sub-systems using physics and data driven models but does not explore the idea of parameter estimation for the overall engine model. Thus, there exists a research opportunity to integrate physics-based models with data-driven approaches for creating efficient surrogates that are suitable for field application.
@@
As a step towards parameter-based health monitoring, we propose an integrated physics and data-based approach to identify unknown parameters of engine sub-systems by using a framework based on physics-informed neural networks [11] (PINNs). In particular, PINNs leverage the known physical laws, typically represented by differential equations, to train neural network parameters for predicting solutions to these equations. Moreover, PINNs can also be applied to predict the unknown parameters of a differential equation system using available output state data. This application mode is termed as the inverse analysis for parameter identification. The nature of these parameters depends on the characteristics of the problem. While PINNs have demonstrated success across a variety of applications [12, 13, 14, 15, 16, 17, 18, 19], their potential for fast prediction in physical systems remains an open research avenue. Thus, our research premise is as follows: given a set of measurements and a physics-based model, estimate the critical parameters for sub-systems that can help in engine health monitoring. In our previous study [20], we demonstrated how PINNs can be used for estimating unknown parameters and gas flow dynamics in a diesel engine model described in [21]. We showed that the proposed method could predict simultaneously both unknown parameters and dynamics accurately. However, a limitation of this work was the high computational cost for parameter estimation, primarily stemming from the need to re-train the PINN model every time inference was required on a new data signal. This limited the applicability of this framework in field systems that require rapid prediction capability. In this study, we propose a hybrid method combining PINNs with neural operator learning for accelerating parameter estimation. While PINNs is employed to estimate the state variable and unknown parameters, the states which are independent are approximated using a neural operator.
@@
Specifically, we consider the Deep Operator Network (DeepONet) as our choice of the neural operator [22] and build on the work of [23], who used DeepONet for predicting the gas flow dynamics described by [21]. This DeepONet model proposed by [23] is not designed for estimating the parameters of critical sub-systems but provides rapid estimates for gas flow dynamics. Hence, our objective is to combine the fast prediction capability of the operator network with the hybrid PINN framework for accelerating parameter estimation on diesel engines. Furthermore, in order to further reduce the computational cost, we present two transfer learning (TL) approaches: a) multi-stage transfer learning, and b) a few-shot transfer learning using a larger multi-head pre-trained neural network. Pre-trained neural networks were considered to approximate empirical formulas in a previous study [20], and the results showed good approximation. One of the major advantages of considering a pre-trained neural network instead of the least squares method is that we do not need to assume the type of non-linearity between the input and the output. The neural network learns the non-linearity itself if sufficient data is provided in the training process. Tosun et al. [3] also observed that a neural network approximation is better than the least squares method. However, one of the major concerns in both neural network and least squares method approximation is how to incorporate the systematic (e.g., manufacturing) variability in parts of the engines in the empirical formulae, i.e., the engine considered to obtain data for the training of the pre-trained network may vary marginally from the engines used in field applications. In this study, we consider dropout in the pre-trained networks to account for such systematic variability.
@@
In summary, we propose a hybrid deep operator-infused Physics-Informed Neural Network (PINN) method to reduce the computation time required for estimating unknown parameters on a diesel engine. The effectiveness and efficiency of the proposed method are validated through several examples. To further reduce the computation cost, we present two transfer learning approaches, a multi-stage transfer learning and a few-shot transfer learning method. The applicability and accuracy of the proposed methods are shown by considering specific examples. For this purpose, we generate training data for the inverse problem, utilizing the Simulink file [24] provided in [21]. We consider the input signal for generating our simulated data using field measurements from engines on a test bed. Additionally, we incorporate appropriate Gaussian noise in the simulated data for the inverse problem to mimic real-world field data and the expected aleatoric uncertainty. Detailed problem statements and different cases considered are discussed in section 1.2 along with a brief discussion on the engine model considered.
1.1 Main contributions
We summarize below the key contribution of the present study aimed at achieving rapid gas flow dynamics prediction and parameter estimation for diesel engine health monitoring.
  • We propose a system-of-systems approach for modeling diesel engines through neural networks by substituting each sub-systems of a diesel engine model with data and physics-based neural networks. Each of these neural networks work in tandem to replicate a complete mean value model of the diesel engine.
@@ 1.1 Main contributions
  • We propose an Operator-infused PINN framework for rapid estimation of unknown system parameters and dynamics in diesel engines. The operator network approximate the independent state variable while PINN is used to model the gas filling dynamics and estimate the unknown parameters.
  • We propose two hybrid transfer learning strategies to meet the computational requirements of field systems. These frameworks utilize transfer learning combined with pre-trained operator networks to meet computational and robustness requirements for field systems.
  • We employ a multi-head training approach utilizing a large corpus of prior training data. This approach also includes a derivative-free layer to calculate physics loss during transfer learning for improved computational efficiency. The multi-head model bears close association to a Foundation model [25] since it can be re-used for different scenarios.
  • We develop an efficient parameter estimation method suited for deployment on field systems. The robustness of our methods was tested by replicating real-world noisy conditions in simulated data.
We organize the rest of the article as follows: Section 2, covers the methodology used to solve the problem stated in section 1.2. We explain the basic concept of our approach in section 2.1 and provide detailed discussion in subsequent subsections. In section 2.3, we present the DeepONet and how we approximate EGR and VGT states. We introduce the hybrid model involving DeepONet and PINN in section 2.4. In section 2.5, we discuss a multi-stage transfer learning strategy followed by a discussion on a computationally efficient few-shot transfer learning strategy in section 2.6. We present the results for Pre-trained networks and DeepONets in section 3. In section 4, we present the results for hybrid-PINN and multi-stage transfer learning. The results of few-shot transfer learning are discussed in section 5. A comparison between the two transfer learning methods is discussed in section 6. Finally, a summary of the present study along with future directions are presented in section 7.
@@ 1.2 Problem setup
A mean value model for the gas flow dynamics of a diesel engine was proposed by [21], and we adopt this model for our present study. The main components of the engine are the intake manifold, the exhaust manifold, the cylinder, the EGR valve system, the compressor, and the turbine. The gas flow dynamics through various components of the engine is described by the state variables of its component and the dependent variables. Specifically, the model we consider [21] has eight states
@@
where $p_{im}$ and $p_{em}$ are the intake and exhaust manifold pressure, respectively, $X_{Oim}$ and $X_{Oem}$ are the oxygen mass fractions in the intake and exhaust manifold, respectively, $\omega_{t}$ is the turbo speed. $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$ are the two states for the EGR actuator dynamics, and $\tilde{u}_{vgt}$ represents the VGT actuator dynamics. The mean value engine model is then expressed as
@@
In this study, the states describing the oxygen mass fraction of the intake and exhaust manifold, i.e., $X_{Oim}$ and $X_{Oem}$, are not considered, as the rest of the states do not depend on these two states. Appendix A provides the governing equations that describe the complete engine model. Interested readers may also refer to [21] or our previous work [20] for more details. The governing equations for the six state variables considered in this work are as follows:
@@
In a real system, measurements corresponding to output states $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ are made along with the control inputs ($\bm{u}$) and engine speed $n_{e}$ through a network of integrated onboard sensors. The intake ($p_{im}$) and exhaust ($p_{em}$) manifold pressures are measured using pressure sensors. The turbocharger shaft speed ($\omega_{t}$) is measured using a rotational speed sensor, while the EGR valve flow rate ($W_{egr}$) is measured using a flow venturi tube. A few critical parameters of the diesel engine that are difficult to estimate on a running engine include the maximum effective area of the EGR valve ($A_{egrmax}$), compensation factor for non-ideal cycles ($\eta_{sc}$), the total heat transfer coefficient of the exhaust pipes ($h_{tot}$), and the maximum area of VGT ($A_{vgtmax}$). By estimating the variation in values of these parameters during engine operation, one can assess the health of the particular sub-system, thereby supporting FDI.
The objectives of the present study are twofold: (1) to predict the gas flow dynamics of all relevant engine states using a physics-informed deep learning framework, and (2) to identify critical unknown parameters in the system, given field measurements of $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$, and $T_{em}$ and system models represented by equation (3)-(10). We refer to this as the inverse parameter identification problem in this work. As discussed in section 1, we aim to formulate a computationally efficient method suitable for use as an on-board monitoring system. The training data required for this study is generated using the Simulink file [24] accompanied with [21]. This Simulink file solves the differential equations discussed above with input vector $\bm{u}$ and engine speed $n_{e}$. Although we consider simulated data for training our neural network model, the input data (i.e. $\bm{u}$ and $n_{e}$) used to generate the simulation results are collected from actual engine test bed, thus representing real-world conditions. Additionally, we assess the effect of aleatoric uncertainty on our model’s performance using Gaussian noise. Thus, we consider two cases for our study:

### v2

# A Digital Twin for Diesel Engines: Operator-infused Physics-Informed Neural Networks with Transfer Learning for Engine Health Monitoring

@@ Abstract
Improving diesel engine efficiency, reducing emissions, and enabling robust health monitoring have been critical research topics in engine modelling. While recent advancements in the use of neural networks for system monitoring have shown promising results, such methods often focus on component-level analysis, lack generalizability, and physical interpretability. In this study, we propose a novel hybrid framework that combines physics-informed neural networks (PINNs) with deep operator networks (DeepONet) to enable accurate and computationally efficient parameter identification in mean-value diesel engine models. Our method leverages physics-based system knowledge in combination with data-driven training of neural networks to enhance model applicability. Incorporating offline-trained DeepONets to predict actuator dynamics significantly lowers the online computation cost when compared to the existing PINN framework. To address the re-training burden typical of PINNs under varying input conditions, we propose two transfer learning (TL) strategies: (i) a multi-stage TL scheme offering better runtime efficiency than full online training of the PINN model and (ii) a few-shot TL scheme that freezes a shared multi-head network body and computes physics-based derivatives required for model training outside the training loop. The second strategy offers a computationally inexpensive and physics-based approach for predicting engine dynamics and parameter identification, offering computational efficiency over the existing PINN framework. Compared to existing health monitoring methods, our framework combines the interpretability of physics-based models with the flexibility of deep learning, offering substantial gains in generalization, accuracy, and deployment efficiency for diesel engine diagnostics.
Keywords Deep Operator Network $\cdot$ Inverse problem $\cdot$ Transfer learning $\cdot$ Few shot learning $\cdot$ Digital twin $\cdot$ System-of-Systems
@@ 1 Introduction
Model-based health monitoring of internal combustion engines has seen significant progress in recent decades [1, 2, 3, 4, 5], motivated by the need to enhance engine performance, reliability, and maintenance efficacy. Engine manufacturers have developed several diagnostic and service tools for operational needs. However, conventional monitoring strategies face persistent challenges, including poor fault isolation, limited prognostic capabilities, a lack of robustness under varying operating conditions, and gradual degradation in diagnostic performance over time. For example, system-level monitors can falsely flag faults due to marginally degraded components acting in combination, leading to expensive trial-and-error approaches for fault identification. To address these limitations, manufacturers have explored component-level monitoring with tight thresholds. This introduces a high rate of false positives and requires the development of high-fidelity models for individual components, resulting in a significant increase in monetary investment and cycle time. Consequently, various alternative solutions have been explored, including the use of complex logic algorithms, multiple degradation thresholds, and out-of-mission system tests. These approaches have led to increasingly sophisticated and expensive diagnostic infrastructures, burdening system design, tuning, and validation processes.
Recent advances in sensing and control technologies have enabled real-time collection of high-fidelity data from the engine’s sub-systems. This has accelerated research interest in the development of tools and methods for engine modelling and diagnostics. Traditional methods such as Extended Kalman Filter (EKF) [6] have been proposed for fault signal correction and diagnosis in the engine’s air intake system. Although widely used, such methods are limited in their ability to capture strong non-linear, transient behaviour observed under real-world operating conditions. To overcome such challenges, numerous studies have explored the use of neural networks for dynamic modelling, fault detection, and virtual sensing in diesel engines. For instance, Biao et al. [7] utilized an auto-regressive neural network framework using engine load, fuel injection, and feedback rotational speed to predict rotational speed and power generation of a sixteen-cylinder locomotive engine with good accuracy. Tosun et al. [8] compared linear regression and neural network for diesel engines fuelled with biodiesel-alcohol mixture to predict engine torque, CO and NOx, illustrating superior results with neural networks. Thompson et al. [9] proposed a neural network framework for heavy-duty diesel engine emission and performance prediction using a data-driven approach. They used a partial recurrent network architecture utilizing instantaneous engine parameter values and 5-10 sec history window of same data as inputs for predicting engine emission and torque prediction within 5% error on field data. Similarly, Wang et al. [10] proposed a neural network method for a virtual NOx sensor aimed at replacing the physical sensors. Their approach utilizes a physical model for generating signals such as EGR flow rate, coolant temperature and others that are then used as an input to a neural network model that predicts NOx levels at transient operating conditions. Shamekhi and Shamekhi [11] proposed Neuro-Mean Value Model that improved the mean value diesel engine model by embedding small, pre-trained neural networks in engine sub-systems.
Building on the advances in applying neural networks for modelling dynamic engine states, efforts have also been made to utilize deep learning methods for fault detection. In an early work, Scaife et al. [12] proposed a single-layer neural network for predicting six output conditions representing the engine’s state of health. This framework was trained using simulated data for healthy and faulty engines, where the fault was simulated with a leak in the exhaust manifold to analyze turbocharger performance degradation. In more recent works, Cervantes-Bobadilla et al. [13] developed a multi-neural network scheme that uses five parallel networks that simultaneously estimate the response from three critical sensors and generate a residual signal by comparing with individual sensor signals during engine operation. The disparity between the physical sensor and neural network estimation was used to trigger a fault signal when a certain threshold is reached. Shahbaz and Amin [14] embedded neural observers within the air-fuel ratio control system for analytical redundancy. In the event of a faulty sensor detected by the engine control unit, the analytical signal generated by the neural observer replaces the physical sensor signal to ensure adequate air-fuel ratio is maintained. Yang et al. [15] proposed a fault detection framework using 1D CNN layers to classify three different fault conditions based on the engine’s vibration signature. The fault conditions were simulated on an engine test bed and the vibration signal was measured using dedicated accelerometers. Wang et al. [16] utilized an auto-encoder setup to identify faulty signals based on the magnitude of reconstruction error between healthy and faulty input signals.
The idea of developing Digital Twins (DT) coupled with deep learning for modelling complex engine systems has also gained popularity. Tsitsilonis et al. [17] proposed a comprehensive DT framework combining a thermodynamic model with operational engine data for component health assessment. This framework successfully identified underperforming cylinders through deviations in key indicators such as Brake Specific Fuel Consumption (BSFC), Indicated Mean Effective Pressure (IMEP), and Exhaust Gas Temperature (EGT). However, their approach relies heavily on first-principles modelling and calibration using controlled test data. Jiang et al. [18] proposed a DT-based auxilliary model with adaptive sparse attention networks to identify valve clearance faults by dynamically filtering redundant features and identifying salient segments of engine vibration data. Khawale et al. [19] extended the DT paradigm by using it for fault identification and active control strategy in their hybrid engine model that incorporates neural networks with conventional engine models. This framework demonstrated the ability to identify faults in injector and manifold leakage, and successfully modify engine parameters to help recover engine performance.
Despite these advances in engine modelling and fault diagnosis, some key limitations persist in literature. The majority of the approaches are purely data-driven which limits interpretability and physical consistency when extrapolating across operating regimes other than what exists in the training set. Additionally, the accuracy of such data-driven approaches rely on the quality and quantity of data that can be collected. To address this, hybrid modelling approaches have emerged with the objective of combining the capability of neural networks with physics-based models. For example, Pulpeiro González et al. [20] proposed a hybrid method combining physics and data to model gas exchange process within diesel engine sub-systems. Although promising, their work is focused on gas flow dynamics of individual components and does not address the dynamics at the system level. In a recent work, Singh et al. [21] utilize a combination of neural network surrogates and physics-based mean value engine model to predict the fault category of diesel engine based. Specifically, they use an auto-encoder to compress an 89-dimensional input signal collected from an engine test bed, combine it with six simulated signals generated by their physics-based model, and use it as an input to a CNN network that classifies the type of engine fault. This method, however depends on identifying deviations between healthy and faulty signals over a period of time which can lead to false prognosis.
A review of existing works, including those using neural networks, indicates a focus on state detection (eg. fault/no-fault classification of known faults) rather than parameter estimation to identify and quantify deviant behaviour. Additionally, typical neural-network based approaches rely purely on empirical data and often generalize poorly outside the training regime. Also, we note that most existing methods focus on component-level modelling and state estimation, instead of complete engine system due to significantly greater complexity associated with such system-wide integration. In our view, monitoring parameter drifts allows estimation of underlying system-parameters based on known physical relationships between components, thus offering a more precise root cause analysis. Also, incorporating parameter identification in health monitoring system helps improve generalization and reduce ambiguity, since these methods are grounded in physics and are more likely to disentangle overlapping effects. In light of these research gaps, this study aims to streamline the development of health monitoring systems, improve fault isolation, system health prognostics, and predictive maintenance capabilities. Our focus is on advancing system-level monitoring solutions that utilize all available data to provide high-fidelity monitoring of system performance while also delivering robust fault isolation and prognostic capabilities.
@@
As a step towards parameter-based health monitoring, we propose an integrated physics and data-based approach to identify unknown parameters of engine sub-systems by using a framework based on physics-informed neural networks (PINNs). In particular, PINNs [22] leverage the known physical laws, typically represented by differential equations, to train neural network parameters for predicting solutions to these equations or predicting unknown parameters of these equations from known data. While PINNs have demonstrated success across a variety of applications [23, 24, 25, 26, 27, 28, 29], their potential for fast prediction in physical systems remains an open research challenge. In a previous study by Nath et al. [30], it was demonstrated how PINNs can be used for estimating unknown parameters and gas flow dynamics in a diesel engine model described in [31]. It was shown that the proposed method could predict simultaneously both unknown parameters and dynamics accurately. However, a limitation of this work was the high computational cost for parameter estimation, primarily stemming from the need to re-train the PINN model every time inference was required on a new data signal. This limited the applicability of this framework in field systems that require rapid prediction capability. Thus, our research premise is as follows: given a set of measurements and a physics-based model, estimate the critical parameters for sub-systems that can help in engine health monitoring. In this study, first, we propose a hybrid method combining PINNs with neural operator learning for accelerating parameter estimation. While PINNs are employed to estimate the state variable and unknown parameters, independent states are approximated using a neural operator. Further, we propose two transfer learning (TL) strategies to reduce the computational cost for application in the real-world.
@@
We consider the Deep Operator Network (DeepONet) [32] as our choice of the neural operator and build on the work of [33], who used DeepONet for predicting the gas flow dynamics described by [31]. This DeepONet model proposed by [33] is not designed for estimating the parameters of critical sub-systems but provides rapid estimates for gas flow dynamics. Hence, our objective is to combine the fast prediction capability of the operator network with the hybrid PINN framework for accelerating parameter estimation on diesel engines. In order to further reduce the computational cost, we present two transfer learning (TL) approaches: a) multi-stage transfer learning and b) a few-shot transfer learning using a larger multi-head pre-trained neural network. In a previous study [30], pre-trained neural networks were used for approximating empirical relationships, and the results showed good matches with the ground truths. One of the major advantages of using a pre-trained neural network over least square-based regression is the fact that neural networks can inherently learn the non-linear mapping between inputs and outputs without assuming the type of relationship. Tosun et al. [8] also observed that a neural network approximation is better than the least squares method. However, one of the major concerns in both neural network and least squares method approximation is how to incorporate the engine’s systematic (e.g., manufacturing) variability in empirical formulae. In other words, the engine unit used to obtain data for training the pre-trained network may vary marginally from the actual engines used in the field. In this study, we consider dropout in the pre-trained networks to account for such systematic variability.
@@
In summary, we propose a hybrid deep operator-infused Physics-Informed Neural Network (PINN) method to reduce the computation time required for estimating unknown parameters on a diesel engine. To further reduce the computation cost, we present two TL approaches, a multi-stage TL and a few-shot TL method. The applicability and accuracy of the proposed methods are shown by considering specific examples for each approach. For this work, we generate training data utilizing the Simulink file [34] provided in [31]. The input signals for generating our simulated data are obtained from field measurements on engine test beds. Additionally, we incorporate appropriate Gaussian noise in the simulated data for the inverse problem to mimic real-world field data and the expected aleatoric uncertainty. Detailed problem statements and different cases considered are discussed in section 1.2 along with a brief discussion on the engine model considered in this study.
1.1 Key research gaps and our contributions
We summarize below the key research gaps we identified and our contributions to address these gaps.
  • Research gap: There is a need to integrate physics-based models with existing data-driven neural network methods for health monitoring in diesel engines. Current neural network methods rely on residual-signal monitoring, making them ambiguous and less objective.
@@ 1.1 Key research gaps and our contributions
    Contribution: We propose an operator-infused PINN framework for rapid estimation of unknown system parameters and dynamics in diesel engines. The operator networks approximate the independent state variables using data-driven training while PINN is used to model the gas filling dynamics and estimate the unknown parameters using known physics from the mean value model. This approach to parameter identification through inverse analysis reduces ambiguity in fault detection, thereby enabling faster decision making.
  • Research gap: Current methods rely on state estimates of sub-systems for health monitoring and fault detection but do not address the engine system holistically.
    Contribution: We propose a system-of-systems approach for modelling diesel engines through neural networks by substituting each sub-system of a diesel engine model with data and physics-based neural networks. Each network works in tandem to replicate a complete mean value model of the diesel engine.
  • Research gap: Real-world application requires methods that can estimate engine dynamics efficiently across different operating conditions. Physics-based methods such as [30] can be extended across different operating scenarios but fail to meet efficiency requirements for field applications.
    Contribution: We develop an efficient parameter estimation method suited for deployment on field systems, which is grounded in physics based knowledge. The robustness of our method is tested by replicating real-world noisy conditions in simulated data. We propose two transfer learning strategies to meet the computational requirements of field systems. These frameworks utilize transfer learning combined with pre-trained operator networks to meet computational and robustness requirements for field systems. Additionally, we employ a multi-head training approach utilizing a large corpus of prior training data. In this method, only a new output layer is trained for a new scenario and the derivatives for the physics loss is calculated only once outside the training loop, further reducing computational complexity. The multi-head model bears close association to a Foundation model [35] since it can be reused for different scenarios or engine configurations.
@@ 1.2 Problem setup
A mean value model for the gas flow dynamics of a diesel engine was proposed by [31], and we adopt this model in our present study. The main components of the engine considered in this model are the intake manifold, the exhaust manifold, the cylinder, the EGR valve system, the compressor, and the turbine. The gas flow dynamics through various components of the engine is described by the state variables of its component and the dependent variables. Specifically, the model we consider [31] has eight states
@@
where $p_{im}$ and $p_{em}$ are the intake and exhaust manifold pressure respectively, $X_{Oim}$ and $X_{Oem}$ are the oxygen mass fractions in the intake and exhaust manifold respectively, $\omega_{t}$ is the turbo speed. $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$ are the two states for the EGR actuator dynamics, and $\tilde{u}_{vgt}$ represents the VGT actuator dynamics. The mean value engine model is then expressed as
@@
In this study, the states describing the oxygen mass fraction of the intake and exhaust manifold, i.e., $X_{Oim}$ and $X_{Oem}$, are not considered, since the rest of the states do not depend on these two states. Appendix section A provides the governing equations that describe the complete engine model. Interested readers may also refer to [31] and [30] for more details. The governing equations for the six state variables considered in this work are as follows:
@@
In a real system, measurements corresponding to output states $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$ and $T_{em}$ are made along with the control inputs ($\bm{u}$) and engine speed $n_{e}$ through a network of integrated onboard sensors. The intake ($p_{im}$) and exhaust ($p_{em}$) manifold pressures are measured using pressure sensors. The turbocharger shaft speed ($\omega_{t}$) is measured using a rotational speed sensor, while the EGR valve flow rate ($W_{egr}$) is measured using a flow venturi tube. $T_{em}$ may be measured using a thermocouple depending on the engine configuration. Some critical parameters of the diesel engine that are difficult to estimate on a running engine include the maximum effective area of the EGR valve ($A_{egrmax}$), compensation factor for non-ideal cycles ($\eta_{sc}$), the total heat transfer coefficient of the exhaust pipes ($h_{tot}$), and the maximum area of VGT ($A_{vgtmax}$). By estimating the variation in values of these parameters during engine operation, one can assess the health of the particular sub-system and identify potential sources of performance issues.
The objectives of the present study are twofold: (1) to predict the gas flow dynamics of all relevant engine states using a physics-informed deep learning framework, and (2) to identify critical unknown parameters in the system, given field measurements of $p_{im}$, $p_{em}$, $\omega_{t}$, $W_{egr}$, and $T_{em}$ and system models represented by equation (3)-(10). We refer to this as the inverse parameter identification problem in this work. As discussed earlier, we aim to formulate a computationally efficient method suitable for use as engine health monitoring system. The training data required for this study is generated using the Simulink file [34] accompanied with [31]. This Simulink file solves the differential equations discussed above with input vector $\bm{u}$ and engine speed $n_{e}$. Although we consider simulated data for training our neural network model, the input data (i.e. $\bm{u}$ and $n_{e}$) used to generate the simulation results are collected from actual engine test bed, thus representing real-world conditions. Additionally, we assess the effect of aleatoric uncertainty on our model’s performance using Gaussian noise. Thus, we consider two cases for our study: